# A candidate quadruple AGN system at

### v1

@@ 4.2 Evolutionary scenario
4.3 Incidence rate of AGN tuples
Following Hennawi et al. (2015), an estimate of the probability to find such two close companion AGNs to a known quasar can be gained from the two-point correlation function for quasars and amounts to , or if considering any halo-scale triple AGN (Appendix D). A different generalized estimate for the AGN tuple incidence rate can be gained by considering the full observed parent sample. Including the candidate presented in this work, 2 AGN tuples have been identified within the QSO MUSEUM survey encompassing 134 quasars at observed with MUSE to similar depths (Arrigoni Battaia et al. 2019; Herwig et al. 2024; González Lobos et al. in preparation). This corresponds to an incidence rate of such systems of 1.5 % with Poisson errors for small-numbers statistics from Gehrels (1986), much larger than the above calculations and than previously believed (Hennawi et al., 2015). This implies undiscovered tuple systems around SDSS quasars (Pâris et al., 2018) in the same redshift range (with a median redshift of ), and, extrapolated to the entire sky, AGN tuples. Assuming that black hole tuples populate the massive end of the halo mass function and that 65 % of them are visible as AGN tuples (Hoffman et al., 2023), this number would amount to all halos down to M⊙ of the halo mass function (Behroozi et al., 2013).
@@ Acknowledgements.
We thank Guinevere Kauffmann for providing useful comments to this work. E.P.F. is supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America.References
@@
- Herwig et al. (2024) Herwig, E., Arrigoni Battaia, F., González Lobos, J., et al. 2024, arXiv e-prints, arXiv:2408.16826
@@ Appendix A Rejecting the possibility of gravitational lensing
In this case, we expect to find a massive object capable of acting as the gravitational lens in between the two images. To calculate the necessary magnitude of a lensing elliptical galaxy, we follow Farina et al. (2013), making use of the Faber–Jackson relation by Nigoche-Netro et al. (2010) and assuming an isothermal sphere to link mass and velocity dispersion. Assuming an Einstein ring radius of , corresponding to half of the angular distance between QSO1 and AGNc1, the minimal possible apparent -band magnitude of an interloper galaxy is at , brighter than both the AGN candidates and QSO1.
We then inspect the residuals of the point spread function modeling (Fig. 1, right). There, an artifact of the CCD gap becomes apparent as vertical line, but no foreground galaxy is revealed. The -band magnitude in a 1.6 aperture between QSO1 and AGNc1 (indicated as pink circle in Fig. 1, right) after masking negative values potentially arising due to the model subtraction is 26.6 mag, far below the faintest possible lensing galaxy. Therefore we conclude that no lensing geometry fits the observed pattern without revealing the lensing galaxy in the data cube.
@@ Appendix B Spectral fitting
We attempt to fit a set of rest-frame UV lines in the spectra of all four sources, namely Ly, N v , C iv , He ii and C iii]. Due to the stark difference in magnitude and therefore data quality, we adopt two different methods. The two SDSS detected quasars QSO1 and QSO2 have sufficient S/N to be fitted using the package pyQSOFIT (Guo et al. 2018; Shen et al. 2019). We use the spectrum extracted from the MUSE cube as explained in Sec. 3.1 for QSO1 as it has higher S/N than the SDSS spectrum. For QSO2, only the SDSS spectrum is available. We smooth both by the same kernel size of pixels. Errors are estimated through 250 MCMC samplings, and the line width is determined from the fit to the C iii]line. Although we attempt to include a component for the N v line when fitting the Ly line complex of QSO2, we do not find a combination of priors that leads to convergence and we thus do not report a line flux for it. From pyQSOFIT we also obtain the continuum luminosity and the FWHM of C iv and use them to calculate black hole mass estimates for QSO1 and QSO2 using the relation and absolute accuracy estimate derived in Vestergaard & Peterson (2006). The obtained values are and .
Due to the low S/N in the continuum, the line spectra of candidates AGNc1 (smoothed by a 1D Gaussian kernel with pixels; Fig. 4) and AGNc2 (smoothed with pixels; Fig. 5) instead are fitted with Gaussian or double Gaussian line models and weighted by the noise spectra extracted from the rescaled variance cube. Fitting is performed using the function curve_fit in scipy (Virtanen et al. 2020).
Due to the highly non-Gaussian Ly line shape of AGNc1, the spectrum does not appear well-fit by the model. To test if this impacts the line flux, we smooth the spectrum further with a Gaussian kernel size of 10 pixels and repeat fitting. We find that the resulting line flux in both Ly and N v is decreases by less than 4 %. For the AGN candidates, the line widths are determined by calculating the second moment in an unsmoothed spectrum after subtracting the mean continuum level. The calculation window is chosen according to the width of the Gaussian line model fitted to the spectrum and centered on the Ly line peak (AGNc1) and the C iv line peak (AGNc2). Although these lines are resonant and therefore not ideal for line width determination, they have the highest S/N and are the only lines allowing for robust calculation of the second moment. Errors on are obtained by resampling the spectra with the associated noise spectra 10,000 times and, after calculating the second moment, the 25th and 75th percentiles of the obtained distribution are taken as lower and upper error.
A major source of uncertainty in the line flux of AGNc1 and AGNc2 is the poorly constrained underlying continuum. Especially the broad lines in the spectrum of AGNc1 are difficult to distinguish from the continuum and are affected by noise fluctuations, necessitating the large smoothing kernel size to be able to fit Gaussian profiles to the lines. Still, we can only place upper limits on C iv in the spectrum of AGNc1, and we also do not detect N v in the spectrum of AGNc2. Considering the large scatter in continuum level, we determine upper limits for the line fluxes by subtracting continuum implied by the minima closest to the expected line peak in the smoothed spectra and calculate moment 0 within 1 left and right of the expected line peak.
@@ Appendix C Estimating the efficiency of ram pressure stripping
The peculiar extended Ly emission associated with the close triple system might be a further indication of on-going galaxy interaction as the emission could be a result of ram-pressure stripped ISM of AGNc2 appearing Ly bright. This could further explain the large velocity differences in the nebula obtained from the first moment with respect to AGNc2 (Fig. 6), reaching 1500 km s-1 at the East edge of the isophote furthest away from the AGN candidate and generally approaching the systemic redshift of AGNc2 closer to it, although the nebula shows significant turbulence. To determine the feasibility of this hypothesis, we estimate the efficiency of stripping in the configuration of the system.
@@
The gravitational pull of AGNc2 withstanding the ram pressure is described by the stellar and gas surface densities and . AGN hosts at typically have a massive (), compact stellar component described by a Sersic profile with an average Sersic index of 2.6 and an effective radius of 0.9 kpc (Kocevski et al. 2023). The gas disk is well-described by an exponential profile, and we assume an effective radius of 2 kpc (Ward et al. 2024) and a cold gas mass fraction of 20 %. The stripped gas mass can then be estimated by integrating the gas mass profile beyond the radius at which ram pressure exceeds the gravitational pull of the galaxy, and it amounts to at least when assuming the line-of-sight velocity, or roughly when using the 3D velocity.
To contrast this with the Ly-bright gas, we can estimate the mass from the extended emission using the equation with the Ly nebula area kpc2 (Herwig et al. 2024), the proton mass and hydrogen fraction and the typical CGM column density of cool gas, log( (Lau et al. 2016). The latter quantity is associated with the highest uncertainty as it can vary by one order of magnitude and it might not be applicable to extraordinary systems. With these values, we obtain an estimated cool gas mass in the Ly nebula of .
Appendix D Probability of finding a similarly close AGN triple
@@
We infer from the quasar luminosity function at (Shen et al. 2020) with a lower magnitude limit of mag and the quasar bolometric luminosity correction from Runnoe et al. (2012) resulting in cMpc-3. Due to the imprecision in redshift determinations from broad QSO emission lines and the possibility of high peculiar velocities causing additional Doppler shifts, we extrapolate the maximum distance in the triplet, , as the highest projected distance plus a 50 % margin to account for projection effects. This yields a quasar-filled comoving volume of cMpc3. On small scales, the two-point correlation function is well described with with and cMpc (Kayo & Oguri 2012). Plugging in these values, we obtain .
@@
The candidate triple AGN presented in this work is closer than previously found AGN triples. To get an estimate of the probability to find any AGN triple on halo scales, we repeat the calculation using kpc, corresponding roughly to the virial radius of a halo with mass , and obtain a probability of . In contrast, the probability to find three AGNs with kpc at random without considering clustering can be described with (Hennawi et al. 2015) and would be exceedingly rare with .

### v2

@@ 4.2 Evolutionary scenario
4.3 Incidence rate of AGN multiplets
Following Hennawi et al. (2015), an estimate of the probability to find such two close companion AGNs to a known quasar can be gained from the two-point correlation function for quasars and, for a magnitude limit of 25 mag, amounts to , or if considering any halo-scale triple AGN (Appendix D). A different generalized estimate for the AGN multiplet incidence rate can be gained by considering the full observed parent sample. Including the candidate presented in this work, 2 AGN multiplets have been identified within the QSO MUSEUM survey encompassing 134 quasars at observed with MUSE to similar depths (Arrigoni Battaia et al. 2019; Herwig et al. 2024; González Lobos et al. in preparation). This corresponds to an incidence rate of such systems of 1.5 % with Poisson errors for small-numbers statistics from Gehrels (1986), much larger than the above calculations and than previously believed (Hennawi et al., 2015). This implies undiscovered multiplet systems around SDSS quasars (Pâris et al., 2018) in the same redshift range (with a median redshift of ), and, extrapolated to the entire sky, AGN multiplet. Assuming that black hole multiplets populate the massive end of the halo mass function and that 65 % of them are visible as AGN multiplets (Hoffman et al., 2023), this number would amount to all halos down to M⊙ of the halo mass function (Behroozi et al., 2013).
@@ Acknowledgements.
We thank the anonymous referee for their constructive comments that helped improve the manuscript and Guinevere Kauffmann for providing useful comments to an earlier version of this work. E.P.F. is supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America.References
@@
- Herwig et al. (2024) Herwig, E., Arrigoni Battaia, F., González Lobos, J., et al. 2024, A&A, 691, A210
@@
- Humphrey et al. (2008) Humphrey, A., Villar-Martín, M., Vernet, J., et al. 2008, MNRAS, 383, 11
@@ Appendix A Rejecting the possibility of gravitational lensing
In this case, we expect to find a massive object capable of acting as the gravitational lens in between the two images. To calculate the necessary magnitude of a lensing elliptical galaxy, we followed Farina et al. (2013), making use of the Faber–Jackson relation by Nigoche-Netro et al. (2010) and assuming an isothermal sphere to link mass and velocity dispersion. Assuming an Einstein ring radius of , corresponding to half of the angular distance between QSO1 and AGNc1, the minimal possible apparent -band magnitude of an interloper galaxy is at , brighter than both the AGN candidates and QSO1.
We then inspected the residuals of the point spread function modeling (Fig. 1, right). There, an artifact of the CCD gap becomes apparent as vertical line, but no foreground galaxy is revealed. The -band magnitude in a 1.6 aperture between QSO1 and AGNc1 (indicated as pink circle in Fig. 1, right) after masking negative values potentially arising due to the model subtraction is 26.4 mag, far below the faintest possible lensing galaxy. Therefore we conclude that no lensing geometry fits the observed pattern without revealing the lensing galaxy in the data cube.
@@ Appendix B Spectral fitting
We attempted to fit a set of rest-frame UV lines to the spectra of all four sources presented in Fig. 2, namely Ly, N v , Si iv , C iv , He ii and C iii]. Due to the stark difference in magnitude and therefore data quality, we adopted two different methods. All line fluxes, equivalent widths (EWs) and line widths are summarized in Table 2.
[…]
B.1 SDSS QSOs
@@
The two SDSS detected quasars QSO1 and QSO2 have sufficient S/N to be fitted using the package pyQSOFIT (Guo et al. 2018; Shen et al. 2019). We used the spectrum extracted from the MUSE cube as explained in Sec. 3.1 for QSO1 as it has higher S/N than the SDSS spectrum. For QSO2, only the SDSS spectrum is available. Errors were estimated through 250 MCMC samplings, and the line width was determined from the fit to the C iii] line. The Gaussian model used to describe the C iii] line is affected by the Gaussian smoothing kernel applied before fitting, and we thus corrected the values down to reflect the intrinsic line width. Although we attempted to include a component for the N v and Si iv line when fitting QSO2, we did not find a combination of priors that leads to convergence and we thus do not report fluxes for these lines. From pyQSOFIT we also obtained the continuum luminosity and the FWHM of C iv and used them to calculate black hole mass estimates for QSO1 and QSO2 using the relation and absolute accuracy estimate derived in Vestergaard & Peterson (2006). With erg s-1, erg s-1, km s-1 and km s-1, we obtained the values and .
B.2 AGN candidates
Due to low S/N, the spectra of candidates AGNc1 and AGNc2 cannot be fitted with pyQSOFIT and we instead subtracted a constant local continuum value before fitting emission lines using the function curve_fit in scipy (Virtanen et al. 2020).
In particular, we obtained a first estimate of the line widths by calculating the second moment of the Ly line (AGNc1) and C iv line (AGNc2) in an unsmoothed spectrum. We used the resulting values of km s-1 and km s-1 respectively (corresponding to and respectively) to constrain the following fit.
Then, we calculated the continuum level and its variance by masking all shaded gray areas in Fig. 2 as well as marked emission lines within (AGNc1) and (AGNc2) of the systemic wavelength. The chosen masking windows are large enough to ensure no line emission is contaminating the continuum calculation and small enough to retain local continuum around each fitted line. The latter is particularly important for the broad line width of AGNc1. Due to the large velocity shift of C iv in AGNc2, we adjusted the wavelength of the line downward by to ensure appropriate masking and fitting. For each fitted emission line, we then considered the next 40 unmasked wavelength channels (corresponding to 50 Å) left and right of the expected peak and calculated the 25th, 50th, and 75th percentile to get the median continuum value (Cont50) and typical range (Cont25, Cont75). Cont25 is consistent with the noise level for most lines (Fig. 4, 5) and thus represents a conservative lower limit for the continuum level as the AGN candidates are detected in continuum emission (Fig. 1). As the determination of the continuum introduces the largest uncertainty in the fit, in particular for the broad lines of AGNc1, we performed three emission line fittings using each of these continuum values for the subtraction and derived conservative line flux errors from the difference in results.
During line fitting, we assumed a single Gaussian model within the bounds given in Table 3, and weighted the fit with noise spectra extracted from the rescaled variance cube. The Ly and N v line of AGNc1 are likely blended together and are thus fitted together using two Gaussian lines centered on the respective systemic wavelengths. From the Gaussian model of the C iii] line, we obtained the line width of km s-1 (AGNc1) and km s-1 (AGNc2). These values are larger than the calculation of the second moment in unsmoothed spectra, but are consistent with line widths in type 1 AGNs and type 2 AGNs respectively. We again corrected the line width of the Gaussian model by the smoothing kernel.
| Object | Normalization | Peak | |
| AGNc1 | (0, inf) | (, ) | |
| AGNc2 | (0, inf) | (, ) |
If the flux value of an emission line is consistent with zero after subtracting Cont75, we consider it as undetected and only place upper limits on the flux. This is the case for C iv (AGNc1) and N v and Si iv (AGNc2). We derived upper limits from the channel-wise RMS of the spectra scaled to and report the limit at significance.
We derived EWs and uncertainties for the high ionization lines Si iv and C iv from the three different Gaussian line fittings and their respective continuum levels and derived upper limits on EWs by assuming the median continuum level Cont50 and the upper flux limit.
@@ Appendix C Estimating the efficiency of ram pressure stripping
The peculiar extended Ly emission associated with the close triple system might be a further indication of on-going galaxy interaction as the emission could be a result of ram-pressure stripped ISM of AGNc2 appearing Ly bright. This could further explain the large velocity differences in the nebula obtained from the first moment with respect to AGNc2 (Fig. 6), reaching 1500 km s-1 at the East edge of the isophote furthest away from the AGN candidate and generally approaching the systemic redshift of AGNc2 closer to it, although the nebula shows significant turbulence. To determine the feasibility of this hypothesis, we estimated the efficiency of stripping in the configuration of the system.
@@
The gravitational pull of AGNc2 withstanding the ram pressure is described by the stellar and gas surface densities and . AGN hosts at typically have a massive (), compact stellar component described by a Sersic profile with an average Sersic index of 2.6 and an effective radius of 0.9 kpc (Kocevski et al. 2023). The gas disk is well-described by an exponential profile, and we assumed an effective radius of 2 kpc (Ward et al. 2024) and a cold gas mass fraction of 20 %. The stripped gas mass can then be estimated by integrating the gas mass profile beyond the radius at which ram pressure exceeds the gravitational pull of the galaxy, and it amounts to at least when assuming the line-of-sight velocity, or roughly when using the 3D velocity.
To contrast this with the Ly-bright gas, we estimated the mass from the extended emission using the equation with the Ly nebula area kpc2 (Herwig et al. 2024), the proton mass and hydrogen fraction and the typical column density of cool gas in the circumgalactic medium, log( (Lau et al. 2016). The latter quantity is associated with the highest uncertainty as it can vary by one order of magnitude and it might not be applicable to extraordinary systems. With these values, we obtained an estimated cool gas mass in the Ly nebula of .
Appendix D Probability of finding a similarly close AGN triplet
@@
We inferred from the quasar luminosity function at (Shen et al. 2020) with a lower magnitude limit of mag and the quasar bolometric luminosity correction from Runnoe et al. (2012) resulting in cMpc-3. Due to the imprecision in redshift determinations from broad QSO emission lines and the possibility of high peculiar velocities causing additional Doppler shifts, we extrapolated the maximum distance in the triplet, , as the highest projected distance plus a 50 % margin to account for projection effects. This yields a quasar-filled comoving volume of cMpc3. On small scales, the two-point correlation function is well described with with and cMpc (Kayo & Oguri 2012). Plugging in these values, we obtained .
@@
The candidate triple AGN presented in this work is closer than previously found AGN triples. To get an estimate of the probability to find any AGN triple on halo scales, we repeated the calculation using kpc, corresponding roughly to the virial radius of a halo with mass , and obtained a probability of . In contrast, the probability to find three AGNs with kpc at random without considering clustering can be described with (Hennawi et al. 2015) and would be exceedingly rare with .
Appendix E Examining possible sources of spectral contamination
Multiple sources of artifacts and spurious signal can contaminate observational data. Here, we examine a few of these effects and determine the level of contamination introduced.
The first artifact to consider is residual emission from the CCD gap visible in the continuum image after subtraction of QSO1, AGNc1 and AGNc2 (Fig. 1, right panel). In particular, the spectrum of AGNc2 overlaps with the CCD gap. Thus, we extracted a spectrum at the same RA as AGNc2, but close to a flux peak on the CCD gap, and rescaled it to match the mean pixel value at the position of AGNc2 in the residual image. This region is indicated by a white plus in Fig. 1. The spectrum, smoothed with the same Gaussian kernel of size pixels, is displayed in Fig. 7. For reference, we also show the spectrum of AGNc2 as well as the expected wavelengths of emission lines.
While the CCD gap might introduce a noisy contribution to the continuum, it does not show strong emission features coinciding with the fitted emission lines of AGNc2.
Next, we examined the level of background emission in the cube by visually selecting ten regions within the same aperture as the spectra with a radius of 0.8 without continuum emission, taking the mean of the extracted spectra and smoothing it with the Gaussian kernel of size pixels. The obtained spectrum is also displayed in Fig. 7 and represents the typical channel-wise background level in the data. Due to its low level, it is not expected to contribute significantly to the continuum emission in any part of the spectra.
A third source of contamination are strong atmospheric emission lines contributing significant flux peaks for the faint sources considered in this work. Despite careful subtraction of the sky spectrum in two different steps during the data reduction and before combining the individual exposures, residuals of lines can still remain in the combined data (Soto et al. 2016). These residuals appear as very thin (one to three pixel wide) emission lines in the spectrum of continuum sources if they were masked together with source emission lines during sky removal. The spectrum of AGNc2 is specifically affected by this as there are multiple narrow lines visible in the unsmoothed spectrum of the combined cube as well as the individual exposures. In Fig. 8, we show the affected spectral windows together with a sky emission model555Generated with the ”The Cerro Paranal Advanced Sky Model”, https://www.eso.org/observing/etc/doc/skycalc/The_Cerro_Paranal_Advanced_Sky_Model.pdf.
In these unsmoothed spectra, the difference in line width to C iv becomes apparent. Each of the narrow emission features not assignable to a common AGN UV line coincides with the position of a sky emission line within one or two pixels. Therefore, these additional lines are most likely sky subtraction residuals.
Finally, given the small number of exposures obtained, individual datacubes could contribute significantly to emission lines, particularly if the line width is narrow. To exclude the possibility that the narrow lines in the spectrum of AGNc2 are spurious, we excluded in turn one of the three exposures during combination of the final cube and examined the Ly and C iv lines in the different resulting spectra in comparison to the spectrum presented in Fig. 2. As shown in Fig. 9, the line emission persists irrespective of the combination of exposures. We caution that the gap between CCDs, which is masked before combination of the cubes, covers parts of the PSF of AGNc2 in the third exposure. However, when combining only two cubes, masked areas in one data cube are filled in with the values of the other data cube. This is the reason why the seemingly decreased line flux of exposure 3 (Fig. 8) does not have a strong affect on the line flux in the combined cube (Fig. 9).